if the line segment intersepted by the parabola y^2=4ax with the line lx+my+n=0 subtends a right angle at the vertex then

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear manan

this problem can be solve easly by hozinisation methord

y2 -4ax =0

lx+my+n=0   or  -(lx +my)/n  = 1

 

put the value of 1 in the above equation

y2 -4ax{-(lx +my)/n} =0

or ny2 + 4alx2 +4amxy =0

this is the pair of equation ,apply the condition of perpendicularity

 n+ 4al =0

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin